# Verbalized Representation Learning for Interpretable Few-Shot Generalization

### v1

@@ Abstract
Humans recognize objects after observing only a few examples, a remarkable capability enabled by their inherent language understanding of the real-world environment. Developing verbalized and interpretable representation can significantly improve model generalization in low-data settings. In this work, we propose Verbalized Representation Learning (VRL), a novel approach for automatically extracting human-interpretable features for object recognition using few-shot data. Our method uniquely captures inter-class differences and intra-class commonalities in the form of natural language by employing a Vision-Language Model (VLM) to identify key discriminative features between different classes and shared characteristics within the same class. These verbalized features are then mapped to numeric vectors through the VLM. The resulting feature vectors can be further utilized to train and infer with downstream classifiers. Experimental results show that, at the same model scale, VRL achieves a absolute improvement over prior state-of-the-art methods while using less data and a smaller mode. Furthermore, compared to human-labeled attributes, the features learned by VRL exhibit a absolute gain when used for downstream classification tasks. Code is available at: link.
@@ 1 Introduction
Prior works [14, 13] have explored integrating language into visual classifiers with a bottleneck of textual attributes. However, these attributes often require human annotations or predefined vocabularies. Some recent methods [3, 19, 44, 7] have sought to automate this process by training neural estimators to transform the statistical significance of the learned features into natural language or explainable outcomes. Yet, these approaches typically demand abundant data to achieve accurate estimations. Large-scale Vision-and-Language Models (VLMs) like CLIP [34] and GPT-4v [1], which are pre-trained on large-scale multimodal datasets, have enabled researchers [27, 40, 33, 10] to perform zero-shot classification with the generated attribute descriptions. However, these attributes are often ungrounded and rely on prior knowledge in pre-training datasets, resulting in low precision when applied to fine-grained or novel concept recognition.
To address the aforementioned challenges, we propose Verbalized Representation Learning (VRL) for automatic, human-interpretable feature extraction using only few-shot data. It applies to fine-grained and novel objects and could work with local models such as LLaVA [26]. Specifically, inspired by self-supervised representation learning (SSL) in computer vision, such as SimCLR [8], MoCo [17], SwAV [6], SimSiam [9], and BYOL [16], we propose to leverage a VLM to capture the inter-class difference and intra-class commonality and articulate these findings in natural language, as illustrated in Fig. 1. Specifically, we cast the VLM to describe the key difference between two images from different classes, which would preserve the discriminative features and remove the redundant ones, such as the yellow body shared by both species on the left side of Fig. 1 (a). Conversely, we employ the VLM to list the key features that are shared by two images within the same class, extracting features that are robust to intra-class variance, such as the orange facial pattern observed in all of the Thalassoma Pavo. Notably, this process can be applied to any two images, whether from different classes or the same class, which allows us to exponentially scale the number of verbalized features with the image samples, enabling effective generalization even in few-shot settings.
@@
We conduct experiments on iNaturalist [39] and Kiki-Bouba dataset [2], where the former includes objects that only have subtle differences between classes, while the latter contains novel objects that are nearly not present in the web-scale datasets. Compared to previous state-of-the-art (SoTA) baselines using 70B models, we achieve a 24 absolute improvement while using 95 less data and a much smaller model with 7B parameters. Against supervised fine-tuned baselines, we observe a 15 absolute improvement. Lastly, when compared to human annotated attributes, the attributes learned by our VRL demonstrate a 20 absolute gain. These results illustrate that the features extracted by VRL not only offer superior effectiveness but also exhibit strong robustness, providing a generalizable solution when adapting to fine-grained or novel classification tasks with a limited data.
@@ 2 Related Work
Natural Language for Image Classification. A common approach for integrating language into visual classifiers involves creating a concept bottleneck [24], where the model first predicts relevant attributes and subsequently uses these attributes to classify the image. Bottleneck methods have been extensively applied in few-shot or zero-shot classification models [14, 13, 24, 21, 25, 15, 36, 7]. However, these methods typically rely on manually annotated or predefined attributes, which can limit their adaptability to novel classes or fine-grained tasks. Recent advancements in VLMs have enabled researchers to directly sample interpretable descriptive features from these models [27, 40, 10, 33, 41]. This approach benefits from incorporating external knowledge embedded in the pre-trained datasets. However, it often relies heavily on the prior knowledge encoded within these models, which can lead to the generation of ungrounded features and difficulty generalizing when the target data is not well-represented in the training datasets. In contrast, our method directly learns grounded, verbalized features from the visual data, allowing for automatic interpretable feature extraction that remains adaptable even for novel classes.
@@ 3.1 Verbalized Representation Learning
Intuitively, and learns complimentary features. captures the most discriminative features while filtering out redundant ones, such as the yellow body present in both types of fish, leading to less noisy features. While is robust to the intra-class variance where it learns to identify the fish based on the shared features, like the pattern on its face, rather than the number of stripes on its back. In the later experimental section, we will further verify this assumption and show that these two features can yield the best performance when combined. In addition, as discussed in the earlier paragraph, even with only -shot samples, our VRL can sample a diverse set of features from negative and positive pairs, making our method excel under low-resource setting. Moreover, we empirically discover that sampling pairs for both negative and positive samples is sufficient to collect a robust set of features.
@@
To obtain feature vectors that can be later used to build visual classifiers, we transform the verbalized features into numeric vectors with the help of VLMs such as CLIP [34] and LLaVA [26]. Given an image and a set of verbalized features, which are in the form of language descriptions, we employ a VLM to determine whether the image possesses those features, as illustrated in Fig. 2 (b), resulting in a feature embedding . Each dimension of is produced by the VLM, indicating the presence of a certain verbalized feature from and or assessing the degree that the image has a certain feature. Concretely, for generative VLMs like LLaVA, each verbalized feature is mapped to 0 or 1, based on whether the model infers the presence of the feature in the image. For VLMs like CLIP, feature embeddings are derived by calculating the similarity between each verbalized feature and the image using CLIP’s visual encoders. This results in a continuous vector indicating the likelihood that the image contains each feature, with an optional similarity threshold to convert it into binary feature vectors.
@@ 4 Experiments
Our experiments are designed to answer the following research questions: (i) How well does VRL generalize to tasks requiring fine-grained recognition with few-shot data? (ii) Can VRL effectively adapt to tasks involving novel concepts that were not part of the VLM’s pre-training resources, using only few-shot data? (iii) How effective does VRL compare to conventional few-shot adaptation algorithms like LoRA fine-tuning or in-context learning? (iv) Do different types of features extracted by VRL mutually benefit each other, and do the extracted features further improve on top of the other commonly used features? (v) How does the performance of automatically features extracted with VRL compare to human-labeled features?
@@ 4.1 Dataset and Evaluation Protocols
We conduct experiments on two different datasets to validate our method’s effectiveness under different scenarios. All reported numbers are the classification accuracy (%).
@@ 4.2 Baseline Methods and Implementation Details
To demonstrate the effectiveness of our method, we compare it against several baselines, including the CLIP-based approach proposed in [10] and other baseline methods that utilize LLMs, such as CBD [27] and LLM-Mutate [10], which leverages LLM’s pretrained knowledge to generate attributes. Additionally, since our method leverages LLaVA-OneVision [26], one LLaVA variant, to generate verbalized features, we perform quantitative comparisons with common approaches used to adapt LLaVA for downstream tasks, such as LoRA fine-tuning [20] and in-context learning [12]. We describe the details of each baseline method as follows.
@@
LLM-Based Methods.
Classification by Description [27] generates a list of attributes by prompting GPT [30] with the class name of the object. While this allows them to leverage the prior knowledge learned by LLMs, the generated features often lack grounding in the actual images and heavily rely on the pre-trained knowledge of LLMs. LLM-Mutate [10] enhances this approach by using the similarity between generated attributes and images to filter out ungrounded ones. However, this method requires access to a full training dataset to accurately estimate relevant attributes and still relies heavily on the capabilities and learned knowledge of the pre-trained LLMs. In our experiments, we not only report results from the original paper but also evaluate this method using the same limited data as our approach for a fair comparison.
@@ 4.3 Fine-Grained Species Classification
We present the results for fine-grained species classification on the iNaturalist dataset in Table 1. We find that compared to previous state-of-the-art (SoTA) methods, even with less data (10 images per species compared to 200+ images per species), and using a significantly smaller model (7B v.s. 70B parameters), our method can already surpass the previous SoTA by . When we increase the scale of our method to a comparable 72B model, we can further increase the improvements to nearly . Moreover, when baseline methods are limited to the same data availability as our approach, our 7B and 72B models achieve improvements of and , respectively. These results highlight the advantage of using grounded descriptions to capture the subtle visual cues, especially when the model needs to discover visual nuances to discriminate species within the same family.
@@ 4.4 Novel Concept Classification
For novel concept classification, we report the results on the Kiki-Bouba dataset in Table 2. Similar to the performance trend observed in the previous experiment, our method surpasses the previous SoTA by with a smaller model, but this time with less data (10 images per species compared to 800+ images per species). Surprisingly, we observe that increasing the number of model parameters does not lead to better performance on this task. One possible explanation for this is that, rather than relying on prior knowledge learned during pre-training, the model needs to focus on learning new, critical features specific to these novel objects with unfamiliar shapes and patterns. Since the objects in this task are unseen during LLM training, existing models struggle to generate relevant attributes and fail to ground visual differences necessary for distinguishing objects from different classes. This hypothesis is further supported by the larger performance gap between our method and those that merely leverage pre-trained LLMs to generate attributes [27]. On this dataset, our method exceeds their performance by , compared to a advantage on the previous dataset.
@@ 4.5 Additional Analysis
We also discover that ensemble classifiers trained separately on each feature type are more effective than simply concatenating all features to build a single classifier. This is primarily because these features are heterogeneous. For example, and are generally binary, with each element indicating the presence of a specific verbalized feature. Therefore, they tend to perform best when used with logistic regression models. Meanwhile, the features encoded by CLIP are continuous, which makes them more suitable for mapping to class labels through MLP models. Empirically, we demonstrate that the best configuration of ensembling can further improves the performance by comparing to feature concatenation. These experiments showcase that our approach is flexible and can be seamlessly integrated with existing representations to build a robust classification model through ensembling.
@@
To evaluate the performance of automatically extracted features compared to human-labeled features, we conduct experiments on the Kiki-Bouba dataset as it provides human-annotated attributes. From Table 4, we observe that VRL extracted features outperform human-labeled attributes across all configurations. Our best configuration, which combines both difference feature vector and commonality feature vector , surpasses the human-derived features by . This underscores the effectiveness of our automatic feature extraction approach in capturing relevant and distinguishing characteristics for classification tasks. These results also suggest that, in scenarios with limited resources or new visual concepts, we can automate the process of feature extraction, reducing the need for manual human annotation.
@@ 4.6 Qualitative Analysis
Figure 3 presents qualitative examples of what verbalized features are extracted from image pairs. We list some of the features that have top-5 feature importance in our trained logistic regression model. Due to the page limit, we provide more qualitative samples in the appendix. For inter-class differences (Fig. 3 (a)), verbalized features are used to capture distinguishing characteristics between different species. For example, the Lichen (Fungi) category is distinguished based on the structure of the thallus, where one image describes a flat, crust-like thallus, and the other describes a complex, branch-like thallus. Similarly, in the Wild Rye (Grass) category, the differences between a slender stem with a feathery inflorescence and a more robust stem with a dense, elongated seed head are verbalized. For intra-class commonality (Fig. 3 (b)), verbalized features are employed to capture shared traits within the same species. For example, the leaves of Manzanita (Berry) are consistently described as “ovate with a pointed tip and slightly serrated edges,” while the inflorescences of Bulrush (Herb) are uniformly described as having a “delicate, feathery appearance.” These verbalized features are used to guide the decision process by helping the model focus on important distinguishing features or shared traits within each class.
@@ A.2 Time Complexity
Given a classification task with classes and -shot examples per class, we are able to construct and pairs for inter-class and intra-class images, respectively. For example, with 5 classes and 10 images per class, we can sample and distinct pairs for inter-class and intra-class cases. In addition, even with the same image pairs, perform sampling during VLM’s generation can also produce diverse verbalized features. Theoretically, to learn a set of verbalized features, the time complexity is for and for , where represents the number of samples generated by the VLM for the same image pair. While it may sound intimidating, empirically, we find that setting to 1 and perform verbalized representation learning on pairs for both inter and intra cases are sufficient to learn a diverse robust features set. Specifically, for a task involving 5 classes with 10 images per class, this requires only 50 inferences, which can be completed in under 30 seconds. After obtaining the verbalized features, each training image must be mapped into numeric representations based on the learned features and . Given that there are images, and each image is evaluated against descriptions, the computational complexity of this stage is . Empirically, for a 5-class 10-shot task, this can be done within 10 minutes.

### v2

@@ Abstract
Humans recognize objects after observing only a few examples, a remarkable capability enabled by their inherent language understanding of the real-world environment. Developing verbalized and interpretable representation can significantly improve model generalization in low-data settings. In this work, we propose Verbalized Representation Learning (VRL), a novel approach for automatically extracting human-interpretable features for object recognition using few-shot data. Our method uniquely captures inter-class differences and intra-class commonalities in the form of natural language by employing a Vision-Language Model (VLM) to identify key discriminative features between different classes and shared characteristics within the same class. These verbalized features are then mapped to numeric vectors through the VLM. The resulting feature vectors can be further utilized to train and infer with downstream classifiers. Experimental results show that, at the same model scale, VRL achieves a absolute improvement over prior state-of-the-art methods while using less data and a smaller model. Furthermore, compared to human-labeled attributes, the features learned by VRL exhibit a absolute gain when used for downstream classification tasks. 111The code is available at: https://github.com/joeyy5588/VRL.
@@ 1 Introduction
Prior works [16, 15] have explored integrating language into visual classifiers with a bottleneck of textual attributes, where the model first predicts relevant attributes and subsequently uses them to classify the image. However, these attributes often require human annotations or predefined ontologies. Development of Large-scale Vision-and-Language Models (VLMs) like CLIP [39] and GPT-4V [1] have enabled recent bottleneck methods [32, 48, 38, 12, 50] to generate relevant attributes by leveraging the prior knowledge of VLMs learned during the pre-training stage. However, these attributes are often ungrounded for the testing images and rely on VLM’s prior knowledge, resulting in low precision when applied to fine-grained or novel concept recognition where objects are underrepresented in the pre-training datasets.
To address the aforementioned challenges, we propose Verbalized Representation Learning (VRL) for automatic, human-interpretable feature extraction using only few-shot data. It applies to fine-grained and novel objects and could work with local models such as LLaVA [28]. Specifically, inspired by self-supervised representation learning (SSL) in computer vision, such as SimCLR [8], MoCo [19], SwAV [6], SimSiam [10], BYOL [18], CLIP-Event [29], and SOLO [11], we propose to leverage a VLM to capture the inter-class difference and intra-class commonality and articulate these findings in natural language, as illustrated in Fig. 1. Specifically, we cast the VLM to describe the key difference between two images from different classes, which would preserve the discriminative features and remove the redundant ones, such as the yellow body shared by both species on the left side of Fig. 1 (a). Conversely, we employ the VLM to list the key features that are shared by two images within the same class, extracting features that are robust to intra-class variance, such as the orange facial pattern observed in all of the Thalassoma Pavo. Notably, this process can be applied to any two images, whether from different classes or the same class, which allows us to exponentially scale the number of useful verbalized features with the image samples, enabling effective generalization even in few-shot settings.
@@
We conduct experiments on iNaturalist [47] and Kiki-Bouba dataset [2], where the former includes objects that only have subtle differences between classes, while the latter contains novel objects that are nearly not present in the web-scale datasets. Compared to previous state-of-the-art (SoTA) baselines using 70B models, we achieve a 24 absolute improvement while using 95 less data and a much smaller model with 7B parameters. Against supervised fine-tuned baselines, we observe a 15 absolute improvement. Lastly, when compared to human annotated attributes, the attributes learned by our VRL demonstrate a 20 absolute gain. These results illustrate that the features extracted by VRL not only offer superior effectiveness but also exhibit strong robustness, providing a generalizable solution when adapting to fine-grained or novel classification tasks with limited data.
@@ 2 Related Work
Natural Language for Image Classification. A common approach for integrating language into visual classifiers involves creating a concept bottleneck [26]. Early bottleneck methods have been extensively applied in few-shot or zero-shot classification models [16, 15, 26, 23, 27, 17, 42, 7]. However, these methods typically rely on manually annotated or predefined attributes, which can limit their adaptability to novel classes or fine-grained tasks. Recent advancements in VLMs have enabled researchers to directly sample interpretable descriptive features from these models [32, 48, 50, 12, 38, 49, 45, 51]. This approach benefits from incorporating external knowledge embedded in the pre-trained datasets. However, it often relies heavily on the prior knowledge encoded within these models, which can lead to the generation of ungrounded features and difficulty generalizing when the target data is not well-represented in the training datasets. In contrast, our method directly learns grounded, verbalized features from the visual data, allowing for automatic interpretable feature extraction that remains adaptable even for novel classes.
@@ 3.1 Verbalized Representation Learning
Intuitively, and learn complimentary features. captures the most discriminative features while filtering out redundant ones, such as the yellow body present in both types of fish, leading to less noisy features. While is robust to the intra-class variance where it learns to identify the fish based on the shared features, like the pattern on its face, rather than the number of stripes on its back. In the later experimental section, we will further verify this assumption and show that these two features can yield the best performance when combined. In addition, as discussed in the earlier paragraph, even with only -shot samples, our VRL can sample a diverse set of features from negative and positive pairs, making our method excel under low-resource setting. Moreover, we empirically discover that sampling pairs for both negative and positive samples is sufficient to collect a robust set of features.
@@
To obtain feature vectors that can be later used to build visual classifiers, we transform the verbalized features into numeric vectors with the help of VLMs such as CLIP [39] and LLaVA [28]. Given an image and a set of verbalized features, which are in the form of language descriptions, we employ a VLM to determine whether the image possesses those features, as illustrated in Fig. 2 (b), resulting in a feature embedding , i.e., . Each dimension of is produced by the VLM, indicating the presence of a certain verbalized feature from and or assessing the degree that the image has a certain feature. Concretely, for generative VLMs like LLaVA, each verbalized feature is mapped to 0 or 1, based on whether the model infers the presence of the feature in the image. For VLMs like CLIP, feature embeddings are derived by calculating the similarity between each verbalized feature and the image using CLIP’s visual encoders. This results in a continuous vector indicating the likelihood that the image contains each feature, with an optional similarity threshold to convert it into binary feature vectors.
@@ 4 Experiments
Our experiments are designed to answer the following research questions: (i) How well does VRL generalize to tasks requiring fine-grained recognition with few-shot data? (ii) Can VRL effectively adapt to tasks involving novel concepts that were not part of the VLM’s pre-training resources, using only few-shot data? (iii) How effective is VRL compared to conventional few-shot adaptation algorithms like LoRA fine-tuning or in-context learning? (iv) Do different types of features extracted by VRL mutually benefit each other, and do the extracted features further improve on top of the other commonly used features? (v) How does the performance of automatically features extracted with VRL compare to human-labeled features?
@@ 4.1 Dataset and Evaluation Protocols
We conduct experiments on iNaturalist and Kiki-Bouba to validate our method’s effectiveness under varied scenarios, with all reported numbers representing classification accuracy (%). Although our primary analysis centers on these datasets, additional results for general few-shot classification on mini-ImageNet [41] are provided in the Appendix B.3.
@@ 4.2 Baseline Methods and Implementation Details
To demonstrate the effectiveness of our method, we compare it against several baselines, including the CLIP-based approach proposed in [12] and other baseline methods that utilize LLMs, such as LaBo [50], CBD [32] and LLM-Mutate [12], which leverage LLM’s pre-trained knowledge to generate attributes. Additionally, since our method leverages LLaVA-OneVision [28] to generate verbalized features, we perform quantitative comparisons with common approaches used to adapt LLaVA for downstream tasks, such as LoRA fine-tuning [22] and in-context learning [14]. We describe the details of each baseline method as follows.
@@
Language Bottleneck Methods.
Classification by Description [32], LaBo [50], and LLM-Mutate [12] generate a list of attributes by prompting LLMs like GPT [35] and Llama [46] with the class name of the object. While this allows them to leverage the prior knowledge learned by LLMs, the generated features often lack grounding in the actual images. LaBo and LLM-Mutate alleviate this issue by using submodular optimization and evolutionary algorithm, respectively, to filter out ungrounded attributes. However, these methods still rely heavily on the capabilities and learned knowledge of the pre-trained LLMs. For LaBo, we adapt the official implementation for our datasets; for LLM-Mutate, we not only report the results from the original paper but also evaluate the method using the same limited data as our approach, ensuring a fair comparison.
@@ 4.3 Fine-Grained Species Classification
We present the results for fine-grained species classification on the iNaturalist dataset in Table 1. We find that compared to previous SoTA methods, even with less data (10 images per species compared to 200+ images per species), and using a significantly smaller model (7B v.s. 70B parameters), our method can already surpass the previous SoTA by . When we increase the scale of our method to a comparable 72B model, we can further increase the improvements to nearly . Moreover, when baseline methods are limited to the same data availability as our approach, our 7B and 72B models achieve improvements of and , respectively. These results highlight the advantage of using grounded descriptions to capture the subtle visual cues, especially when the model needs to discover visual nuances to discriminate species within the same family.
@@ 4.4 Novel Concept Classification
For novel concept classification, we report the results on the Kiki-Bouba dataset in Table 2. Similar to the performance trend observed in the previous experiment, our method surpasses the previous SoTA by with a smaller model, but with less data (10 images per species compared to 800+ images per species). Surprisingly, we observe that increasing the number of model parameters does not lead to better performance on this task. One possible explanation for this is that, rather than relying on prior knowledge learned during pre-training, the model needs to focus on learning new, critical features specific to these novel objects with unfamiliar shapes and patterns. Since the objects in this task are unseen during LLM training, existing models struggle to generate relevant attributes and fail to ground visual differences necessary for distinguishing objects from different classes. This hypothesis is further supported by the larger performance gap between our method and those that rely on LLM’s pre-trained knowledge [32]. On this dataset, our method exceeds their performance by , compared to a advantage on the previous dataset.
@@ 4.5 Additional Analysis
We also discover that ensemble classifiers trained separately on each feature type are more effective than simply concatenating all features to build a single classifier. This is primarily because these features are heterogeneous. For example, and are generally binary, with each element indicating the presence of a specific verbalized feature. Therefore, they tend to perform best when used with logistic regression models. Meanwhile, the features encoded by CLIP are continuous, which makes them more suitable for mapping to class labels through MLP models. Empirically, we demonstrate that the best configuration of ensembling can further improves the performance by comparing to feature concatenation. We find that with three classifiers, hard-voting is prone to high variance, while soft-voting is more robust since it leverages the confidence of each classifier by averaging their probabilities. These experiments showcase that our approach is flexible and can be seamlessly integrated with existing representations to build a robust classification model through ensembling.
@@
To evaluate the performance of automatically extracted features compared to human-labeled features, we conduct experiments on the Kiki-Bouba dataset using the human-annotated attributes provided by [12]. From Table 4, we observe that VRL extracted features outperform human-labeled attributes across all configurations. Our best configuration, which combines both difference feature vector and commonality feature vector , surpasses the human-derived features by . This underscores the effectiveness of our automatic feature extraction approach in capturing relevant and distinguishing characteristics for classification tasks. These results also suggest that, in scenarios with limited resources or new visual concepts, we can automate the process of feature extraction, reducing the need for manual human annotation.
@@ 4.6 Qualitative Analysis
Figure 3 presents qualitative examples of what verbalized features are extracted from image pairs. We list some of the features that have top-5 feature importance in our trained logistic regression model. For inter-class differences (Fig. 3 (a)), verbalized features are used to capture distinguishing characteristics between different species. For example, the Lichen (Fungi) category is distinguished based on the structure of the thallus, where one image describes a flat, crust-like thallus, and the other describes a complex, branch-like thallus. Similarly, in the Wild Rye (Grass) category, the differences between a slender stem with a feathery inflorescence and a more robust stem with a dense, elongated seed head are verbalized. For intra-class commonality (Fig. 3 (b)), verbalized features are employed to capture shared traits within the same species. For example, the leaves of Manzanita (Berry) are consistently described as “ovate with a pointed tip and slightly serrated edges,” while the inflorescences of Bulrush (Herb) are uniformly described as having a “delicate, feathery appearance.” These verbalized features are used to guide the decision process by helping the model focus on important distinguishing features or shared traits within each class. Due to page limit, see Appendix B.5 for human study on the learned verbalized features.
@@ 5 Conclusion
Acknowledgements
We thank anonymous reviewers and other members of UCLA-NLP+ group for their helpful comments. This work was partially supported by U.S. DARPA ECOLE Program No. #HR00112390060, ONR grant N00014-23-1-2780, DARPA ANSR program FA8750-23-2-0004, and Apple. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ A.2 Time Complexity
Given a classification task with classes and -shot examples per class, we are able to construct and pairs for inter-class and intra-class images, respectively. For example, with 5 classes and 10 images per class, we can sample and distinct pairs for inter-class and intra-class cases. In addition, even with the same image pairs, perform sampling during VLM’s generation can also produce diverse verbalized features. Theoretically, to learn a set of verbalized features, the time complexity is for and for , where represents the number of samples generated by the VLM for the same image pair. While it may sound intimidating, empirically, we find that setting to 1 and perform verbalized representation learning on pairs for both inter and intra cases are sufficient to learn a diverse robust features set. Specifically, for a task involving 5 classes with 10 images per class, this requires only 50 inferences, which can be completed in under 30 seconds. After obtaining the verbalized features, each training image must be mapped into numeric representations based on the learned features and . Given that there are images, and each image is evaluated against descriptions, the computational complexity of this stage is . Empirically, for a task with 5 classes and 10 images, our approach requires less than 30 seconds on a single A6000 to extract verbalized features. In contrast, LLM-Mutate [12] based on text-based LLM sampling could take 22 hours as the generated features often lack visual grounding, resulting in slower convergence on discriminative features.